OpenStudy (anonymous):

Solve for x. -3 + 8x - 5 = -8

OpenStudy (anonymous):

-3 + 8x - 5 = -8 +3 +5 +3 +5 8x = 0 (8x = 0)/8 x = 0

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (anonymous):

No problem ^_^

OpenStudy (hhopke):

x=0 because... -3 - 5 = -8. (-8+8x=-8) -8+8=0 (8x=0) x=0 / 8x 0 \ | -- = -- = 0 | \ 8 8 /

OpenStudy (hhopke):

I'm a slow typer.

OpenStudy (anonymous):

Yep, we know that ^_^ Thanks for trying, @hhopke

OpenStudy (anonymous):

It's the thought that counts

OpenStudy (anonymous):

-3x - 8 + 6x = 7 that's okay thanks for your help as well

OpenStudy (anonymous):

very creative as well (:

OpenStudy (anonymous):

Thank you Do you need help solving -3x - 8 + 6x = 7?